OpenStudy (anonymous):

Your English teacher has given you a list of 10 possible topics for essays. You need to pick four topics from the list to write about over the school year. How many different groups of four essay topics can you choose?

OpenStudy (goformit100):



OpenStudy (compassionate):

(2^4 = 16)

OpenStudy (lgbasallote):

\[\LARGE ^{10} C _4 \implies \frac{10!}{4! (10-4)!}\]

OpenStudy (lgbasallote):

\[\LARGE \implies \frac{10!}{4! 6!} \implies \frac{ 10 \times 9 \times 8 \times 7 \times \cancel{6!}}{4! \cancel{6!}}\]

OpenStudy (anonymous):

confuseddd